OpenStudy (anonymous):

how to calculate log(1/0.02) without a calculator.

hero (hero):

I thought this was going to be a tutorial where the OP would explain how to do it.

OpenStudy (anonymous):

\[Log(\frac{a}{b}) = Log(a) - Log(b)\] Use this and try to solve further...

OpenStudy (anonymous):

Yeah honestly the closest you can get is either log(50) or log(1) - log(0.02) = -log(0.02) in either case not much you can do.

hero (hero):

How would you do log (a) without a calculator? lol

OpenStudy (anonymous):

Maybe its not base 10?

hero (hero):

\[\log_{10} x = y \equiv 10^y = x\]

OpenStudy (anonymous):

\[Log(1) - Log(0.02)\] Log1 is 0..

OpenStudy (anonymous):

How will you find y without using calculator??

OpenStudy (anonymous):

@bronzegoddess what is the base of your log?

OpenStudy (anonymous):

\[0 - Log(\frac{100}{2}) = - [Log(100) - Log(2)]\] Log(100) = 2 because: \[Log(100) = Log(10)^2 = 2Log(10) = 2 \times 1 = 2\] \[ = -[2 - Log(2)] = [Log(2) - 2]\] Log(2) has standard value which we should remember Log(2) = 0.3010 Put in that: \[(2 - 0.3010)\] This will be answer...

hero (hero):

I was going to use 100/2

OpenStudy (anonymous):

No, I was

OpenStudy (anonymous):

Then use...

OpenStudy (anonymous):

No, you do it

OpenStudy (anonymous):

the base is 10

OpenStudy (anonymous):

I think when base is not mention, then we take it as 10.. Well, I think so..

OpenStudy (anonymous):